OpenStudy (anonymous):

how do I find the lowest common denominator of fractions, per say this one inparticular ? (3/4-2/3)*(1/2+1/3)

OpenStudy (anonymous):

LCD is 12

OpenStudy (precal):

look at 4 3 2 3 and their factors 4=2*2 3=3*1 2=2*1 3=3*1 take 2*2*3=12